OpenStudy (anonymous):

Choose the value of the y determinant (Dy) in the following system. 2x + 4y = 5 3x – y = 7

OpenStudy (anonymous):

the options are A. –14 B. –33 C. –1 D. 6

OpenStudy (bahrom7893):

What is the equation for the determinant?

OpenStudy (anonymous):

isn't it with cramer's rule?

OpenStudy (anonymous):

i don't know how to plug it in

OpenStudy (bahrom7893):

Well this is going to be your matrix: 2 4 |5 3 -1 |7

OpenStudy (anonymous):

so you just write the value of A,B and C?

OpenStudy (anonymous):

now you substitute the y column right? 2 5 3 7

OpenStudy (bahrom7893):

Give me a moment..

OpenStudy (anonymous):

thank you!! i'll check it out

OpenStudy (bahrom7893):

It's pretty much the same thing, only different numbers.

OpenStudy (anonymous):

so it's -1?

OpenStudy (bahrom7893):

2 5 3 7 2*7 - 5*3 = 14-15=-1 yup

OpenStudy (anonymous):

thanks!!